OpenStudy (anonymous):

ok well the fractions are 2/3 and 3/4 i don't know how to figure out if they are equal or not

hartnn (hartnn):

the 2 fractions a/b and c/d are equal if ad=bc using this, can u find whether 2/3 and 3/4 are equal?

OpenStudy (anonymous):

no i'm confused because a & d are dirfferent does that make it not equal ?

OpenStudy (anonymous):

or let me ask this how i know if it's equal?? what do I do to find that out

hartnn (hartnn):

u need to find the product ad by multiplying, so here what is 2*4 =? then u need to find the product b*c, here 3*3=? if these 2 values are equal, only then the fraction is equal.

OpenStudy (anonymous):

2/3 and 3/4 are not equal, if you multiply 2/3 by 4/4 and 3/4 by 2/3 (so they have the same denominator) you'll see it's 8/12 compared to 6/12 so 2/3 is greater.

OpenStudy (anonymous):

ok thanks so much this really helped :)

OpenStudy (anonymous):

*sorry 3/4 by 3/3 so it equals 9/12 so 3/4 is greater